Simplify the sum Σ from r=0 to 2n+1 of [3/(r+1)(r+2)]


1
Expert's answer
2021-11-22T12:15:08-0500

Decomposing into two fractions



"\frac{3}{(r+1)(r+2)}=\frac{A}{r+1}+\frac{B}{r+2}"


Multiplying both sides by (r+1)(r+2)


"3=A(r+2)+B(r+1)"


Let "r=-2"

"\implies3=B(-2+1)=-B\implies \>B=-3"


Let "r=-1"

"\implies3=A(-1+2)\implies\>A=3"



"\therefore\frac{3}{(r+1)(r+2)}=\frac{3}{r+1}-\frac{3}{r+2}"



Writing the first three and the last there terms


"\displaystyle\sum_{r=0}^{2n+1}(\frac {3}{r+1}-\frac{3}{r+2})=({3-\frac {3}{2}})+(\frac{3}{2}-\frac{3}{3})+(\frac{3}{3}-\frac{3}{4})+...\\\>+(\frac{3}{2n}-\>\>\frac{3}{2n+1})+(\frac{3}{2n+1}-\frac{3}{2n+2})+(\frac{3}{2n+2}-\frac{3}{2n+3})"


All fraction cancels apart from 1st and last


"\implies" Summation "=3-\frac{3}{2n+3}=\frac{6n+9-3}{2n+3}"



"\therefore" "\displaystyle\sum_{r=0}^{2n+1}" "\frac{3}{(r+1)(r+2)}=\frac{6n+6}{2n+3}=\frac{6(n+1)}{2n+3}"